# A conjecture of Radu and Sellers on congruences modulo powers of 2 for broken 3-diamond partitions

### v1

@@ Abstract.
In 2007, Andrews and Paule introduced the family of functions $\Delta_{k}(n)$, which enumerate the number of broken $k$-diamond partitions for a fixed positive integer $k$. In 2013, Radu and Sellers completely characterized the parity of $\Delta_{3}(8n+r)$ for some $r$ and gave a conjecture on congruences modulo powers of 2 for broken 3-diamond partitions. We use an unconventional $U$-sequence to solve the revised conjecture proposed by Radu and Sellers.
@@ 1. Introduction
A partition of a positive integer $n$ is defined as a sequence of positive integers in non-increasing order that sum to $n$. Denoted by $p(n)$, it represents the count of such partitions for $n$. The following distinguished congruences were discovered by Ramanujan in [13]:
@@
for every non-negative integer $n$. In fact, for $\alpha\geq 1$
@@
(1.2) $p(7^{\alpha}n+\delta_{7,\alpha})$ $\equiv 0\pmod{7^{[\frac{\alpha+2}{2}]}},$
@@
where $\delta_{t,\alpha}$ is the reciprocal modulo $t^{\alpha}$ of 24. (1.1) and (1.2) were first proved by Watson [16] in 1938. For an elementary of (1.1) see Hirschhorn and Hunt [8] and for an elementary proof of (1.2) see Garvan [6]. (1.3) was proved by Atkin [3] in 1967.
In 2007, Andrews and Paule [2] introduced the family of functions $\Delta_{k}(n)$ which enumerate the number of broken $k$-diamond partitions for a fixed positive integer $k$. These are constructed in such a way that the generating functions of their counting sequences $\{\Delta_{k}(n)\}_{n\geq 0}$ are closely related to modular forms. Namely,
@@
$\eta(\tau)=q^{\frac{1}{24}}\prod_{n=1}^{\infty}(1-q^{n})~{}~{}(q:=e^{2\pi i\tau}).$
For $k=1$, Andrews and Paule proved that for all $n\geq 0$, $\Delta_{1}(2n+1)\equiv 0\pmod{3}$. Hirschhorn and Sellers [9] considered cases of $k=1$ and $k=2$ and gave the parity results below: for all $n\geq 0$,
@@
Besides, Chan [5] provided a different proof of the parity results for $\Delta_{2}$ mentioned above as well as a number of congruences modulo powers of 5. Paule and Radu [10] also proved a number of congruences modulo 5 for broken 2-diamond partitions, and they also shared conjectures related to broken 3-diamond partitions modulo 7 and broken 5-diamond partitions modulo 11, two of them proved by Xiong [15].
In 2013, Radu and Sellers [12] completely characterized the parity of $\Delta_{3}(8n+r)$ for some $r$ and gave the following conjecture modulo powers of $2$.
@@
They also proved the case for $\alpha=1$. Later, Xia [14] demonstrated that (1.4) holds for $\alpha=2$ and (1.5) holds for all $\alpha\geq 1$. But (1.4) is incorrect in general. For example let $\alpha=3$ and $n=1$, we have
@@
In this paper, we prove this conjecture by following the method in [11]. We associate the family of congruences with a sequence of functions defined on the classical modular curve $X_{0}(28)$ (a curve of genus 2 and cusp count 6). Let
@@
We mention that we use an unconventional $U$-sequence because $F(q)/F(q^{4})$ is not a modular function, and there does not exist any $U$-sequence of functions defined on the classical modular curve $X_{0}(14)$ (a curve of genus 1 and cusp count 4) associated with the family of congruences.
The paper is organized as follows. In Section 2 we introduce the necessary notations and definitions. In Section 3 we derive modular equations on $\Gamma_{0}(28)$. Section 4 is dedicated to the proof of revised Conjecture 1.1.
@@ 2. Preliminaries
In this paper, we use the following conventions: =$\mathbb{N}^{\ast}=\{1,2,\cdots\}$ denotes the set of positive integers. The complex upper half plane is denoted by :=$\mathbb{H}:=\{\tau\in\mathbb{C}:\Im(\tau)\textgreater 0\}$. We will also use the short hand notation for the Dedekind eta function:
$\eta_{n}(\tau):=\eta(n\tau),~{}~{}n\in\mathbb{Z},~{}\tau\in\mathbb{H}.$
@@
Let $f=\sum_{n\in\mathbb{Z}}a_{n}q^{n}$, $f\neq 0$, be such that $a_{n}=0$ for almost all $n\textless 0$. We define the order of $f$ (with respect to $q$) as the smallest integer $N$ such that $a_{N}\neq 0$ and write $N=ord_{q}(f)$. For instance, let $F=f\circ t=\sum_{n\in\mathbb{Z}}a_{n}t^{n}$ where $t=\sum_{n\geq 1}b_{n}q^{n}$, then the $t$-order of $F$ is defined to be the smallest integer $N$ such that $a_{N}\neq 0$ and write $N=ord_{t}(F)$.
@@
For $f:\mathbb{H}\rightarrow\mathbb{C}$ and $m\in\mathbb{N}^{\ast}$ we define $U_{m}(f):\mathbb{H}\rightarrow\mathbb{C}$ by
$U_{m}(f)(\tau):=\frac{1}{m}\sum_{\lambda=0}^{m-1}f(\frac{\tau+\lambda}{m}),~{}\tau\in\mathbb{H}.$
$U_{m}$ is linear (over $\mathbb{C}$); in addition, it is easy to verify that
$U_{mn}=U_{m}\circ U_{n}=U_{n}\circ U_{m},~{}m,n\in\mathbb{N}^{\ast}.$
The operators $U_{m}$, introduced by Atkin and Lehner [4], are closely related to Hecke operators. They typically arise in the context of partition congruences [1] mostly because of the property: if
$f(\tau)=\sum_{n=-\infty}^{\infty}f_{n}q^{n}~{}~{}(q=e^{2\pi i\tau}),$
@@
For $f:\mathbb{H}\rightarrow\mathbb{C}$ we define $U^{(0)}(f),U^{(1)}(f):\mathbb{H}\rightarrow\mathbb{C}$ by $U^{(0)}(f):=U_{2}(Af)$ and $U^{(1)}(f):=U_{2}(Bf)$, where $A:=\frac{\eta_{8}^{4}\eta_{2}^{2}}{\eta_{4}^{2}\eta_{1}^{4}}$ and $B:=\frac{\eta_{1}^{4}\eta_{2}^{2}\eta_{8}^{4}}{\eta_{4}^{10}}$.
@@
$L_{2\alpha-1}:=U^{(0)}(L_{2\alpha-2})~{}~{}~{}~{}~{}\text{and}~{}~{}~{}L_{2\alpha}:=U^{(1)}(L_{2\alpha-1}).$
The proof of the following lemma is completely analogous to [3, p. 23] and we omit it.
@@
$t:=\frac{\eta_{28}\eta_{4}}{\eta_{7}\eta_{1}},~{}~{}~{}p:=\frac{1}{2}(1+\frac{7\eta_{14}^{7}\eta_{4}^{5}\eta_{1}^{6}}{\eta_{28}^{3}\eta_{7}^{2}\eta_{2}^{13}}),$
$p_{0}:=\frac{\eta_{7}\eta_{4}^{7}}{\eta_{28}\eta_{1}^{7}},~{}~{}~{}y_{0}:=\frac{7\eta_{14}^{5}\eta_{4}^{5}\eta_{2}}{\eta_{28}^{3}\eta_{1}^{8}},~{}~{}~{}p_{1}:=\frac{1}{4}(p+2pt+t)+t^{2}.$
@@ 3. Modular equation
Our proof of Radu and Sellers’ conjecture relies on the identities in the Appendix. All these identities can be proved using modular function theory. Regarding to the valence formula in the modular forms, Garvan has written a MAPLE package called ETA which can prove Group I, Group III, and Group IV in the Appendix automatically. See
@@
The tutorial for this see [7]. The identities in Group II of the Appendix cannot be proved directly since $p_{1}$ is not an eta-quotient. However, we can use the ETA MAPLE package to prove that
(3.2) $U^{(0)}(t^{3})=$ $p_{0}(256t^{7}+640t^{6}+864t^{5}+656t^{4}+276t^{3}+64t^{2}+7t)$
$-y_{0}(32t^{5}+48t^{4}+28t^{3}+8t^{2}+t);$
(3.3) $U^{(0)}(t^{4})=$ $p_{0}(1024t^{9}+3072t^{8}+4864t^{7}+4672t^{6}+2832t^{5}+1104t^{4}+260t^{3}+28t^{2})$
$-y_{0}(128t^{7}+256t^{6}+224t^{5}+112t^{4}+32t^{3}+4t^{2});$
@@
where $y:=2p-1$ be an eta-quotient. Then we deduce the identities in Group II using (3.2)–(3.6) along with the identities in Group I since $p_{1}=\frac{1}{8}+\frac{1}{2}t+t^{2}+\frac{1}{8}y+\frac{1}{4}yt$.
@@
With $t=t(\tau)$ as in Definition 2.5 we define:
$a_{0}(t)=2t^{2}+t,~{}~{}~{}a_{1}(t)=-4t^{2}-2t.$
@@
(3.7) $a_{j}(t)=\sum_{l=1}^{2}s(j,l)2^{\lfloor\frac{2l+j-1}{2}\rfloor}t^{l}.$
@@
$t_{\lambda}(\tau):=t(\frac{\tau+\lambda}{2}),~{}~{}\tau\in\mathbb{H}.$
Then in the polynomial ring $\mathbb{C}(t)[X]$:
@@
$a_{0}(t)=$ $t_{0}t_{1}=\prod_{\lambda=0}^{1}q^{1/2}\prod_{n=1}^{\infty}\frac{(1-q^{14n})(1-q^{2n})}{(1-(-1)^{\lambda n}q^{7n/2})(1-(-1)^{\lambda n}q^{n/2})}$
$=$ $q\prod_{n=1}^{\infty}(1-q^{14n})^{2}(1-q^{2n})^{2}\prod_{n=1}^{\infty}\frac{1}{(1-q^{2n+1})(1-q^{14n+7})}\prod_{n=1}^{\infty}\frac{1}{(1-q^{n})^{2}(1-q^{7n})^{2}}$
$=$ $q\prod_{n=1}^{\infty}\frac{(1-q^{14n})^{3}(1-q^{2n})^{3}}{(1-q^{7n})^{3}(1-q^{n})^{3}}=2t^{2}+t,$
where the last equation was verified by Garvan’s MAPLE package ETA (3.1).
@@
where the last equation was again verified by the ETA MAPLE package. ∎
@@
For $u:\mathbb{H}\rightarrow\mathbb{C}$ and $j\in\mathbb{Z}$:
@@
Multiplying both sides with $u_{\lambda}t_{\lambda}^{j-2}$ where $u_{\lambda}(\tau):=u((\tau+\lambda)/2)$ gives
@@
Summing both sides over all $\lambda$ from $\{0,1\}$ completes the proof of the lemma. ∎
According to Lemma 3.3, we can calculate $U^{(0)}(t^{k})$, $U^{(0)}(p_{1}t^{k})$, $U^{(1)}(y_{0}t^{k})$, and $U^{(1)}(p_{0}t^{k})$ using the identities in Appendix for all $k\geq 1$ recursively. For example
(3.8) $U^{(1)}(y_{0}t^{3})=$ $(4t^{2}+2t)U^{(1)}(y_{0}t^{2})-(2t^{2}+t)U^{(1)}(y_{0}t)$
$=$ $p_{1}(2^{17}t^{13}+2^{19}t^{12}+125\cdot 2^{13}t^{11}+315\cdot 2^{12}t^{10}+2261\cdot 2^{9}t^{9}+1505\cdot 2^{9}t^{8}$
$+6031\cdot 2^{6}t^{7}+1131\cdot 2^{7}t^{6}+4953\cdot 2^{3}t^{5}+1875\cdot 2^{2}t^{4}+441\cdot 2t^{3}+49t^{2})$
$-2^{17}t^{15}-9\cdot 2^{16}t^{14}-145\cdot 2^{13}t^{13}-93\cdot 2^{14}t^{12}-2693\cdot 2^{9}t^{11}-3591\cdot 2^{8}t^{10}$
$-7165\cdot 2^{6}t^{9}-5305\cdot 2^{5}t^{8}-5675\cdot 2^{3}t^{7}-1035\cdot 2^{3}t^{6}-463\cdot 2t^{5}-13\cdot 2^{2}t^{4}.$

### v2

@@ Abstract.
In 2007, Andrews and Paule introduced the family of functions $\Delta_{k}(n)$, which enumerate the number of broken $k$-diamond partitions for a fixed positive integer $k$. In 2013, Radu and Sellers completely characterized the parity of $\Delta_{3}(8n+r)$ for certain values of $r$ and proposed a conjecture on congruences modulo powers of $2$ for broken $3$-diamond partitions. In this paper, we employ an unconventional $U$-sequence to resolve the revised conjecture put forward by Radu and Sellers.
@@ 1. Introduction
A partition of a positive integer $n$ is defined as a sequence of positive integers in non-increasing order that sum to $n$. Denoted by $p(n)$, it represents the number of such partitions of $n$. The following celebrated congruences were discovered by Ramanujan in [13]:
@@
for every non-negative integer $n$. In fact, for $\alpha\geq 1$,
@@
(1.2) $p(7^{\alpha}n+\delta_{7,\alpha})$ $\equiv 0\pmod{7^{\lfloor\frac{\alpha+2}{2}\rfloor}}$,
@@
where $\delta_{t,\alpha}$ is the multiplicative inverse of $24$ modulo $t^{\alpha}$. Congruences (1.1) and (1.2) were first proved by Watson [16] in 1938. For an elementary proof of (1.1), see Hirschhorn and Hunt [8]; for an elementary proof of (1.2), see Garvan [6]. Congruence (1.3) was proved by Atkin [3] in 1967.
In 2007, Andrews and Paule [2] introduced the family of functions $\Delta_{k}(n)$, which enumerate the number of broken $k$-diamond partitions for a fixed positive integer $k$. These are constructed such that the generating functions of their counting sequences $\{\Delta_{k}(n)\}_{n\geq 0}$ are closely related to modular forms. Namely,
@@
$\eta(\tau)=q^{\frac{1}{24}}\prod_{n=1}^{\infty}(1-q^{n})\qquad(q:=e^{2\pi i\tau}).$
For $k=1$, Andrews and Paule proved that for all $n\geq 0$, $\Delta_{1}(2n+1)\equiv 0\pmod{3}$. Hirschhorn and Sellers [9] considered the cases $k=1$ and $k=2$ and obtained the following parity results: for all $n\geq 0$,
@@
Besides, Chan [5] provided a different proof of the parity results for $\Delta_{2}$ mentioned above, as well as a number of congruences modulo powers of $5$. Paule and Radu [10] also proved several congruences modulo $5$ for broken $2$-diamond partitions, and they also shared conjectures related to broken $3$-diamond partitions modulo $7$ and broken $5$-diamond partitions modulo $11$, two of which were proved by Xiong [15].
In 2013, Radu and Sellers [12] completely characterized the parity of $\Delta_{3}(8n+r)$ for certain $r$ and proposed the following conjecture modulo powers of $2$.
@@
They also proved the case $\alpha=1$. Later, Xia [14] demonstrated that (1.4) holds for $\alpha=2$ and that (1.5) holds for all $\alpha\geq 1$. However, (1.4) is incorrect in general. For example, let $\alpha=3$ and $n=1$; we have
@@
In this paper, we prove this conjecture by following the method in [11]. We associate the family of congruences with a sequence of functions defined on the classical modular curve $X_{0}(28)$ (a curve of genus $2$ with cusp count $6$). Let
@@
We note that we employ an unconventional $U$-sequence, since $F(q)/F(q^{4})$ is not a modular function, and there is no $U$-sequence of functions defined on the classical modular curve $X_{0}(14)$ (a curve of genus $1$ with cusp count $4$) associated with the family of congruences.
The paper is organized as follows. In Section 2 we introduce the necessary notations and definitions. In Section 3 we derive modular equations on $\Gamma_{0}(28)$. Section 4 is dedicated to the proof of the revised Conjecture 1.1.
@@ 2. Preliminaries
In this paper, we use the following conventions: $\mathbb{N}^{\ast}=\{1,2,\dots\}$ denotes the set of positive integers. The complex upper half-plane is denoted by $\mathbb{H}:=\{\tau\in\mathbb{C}:\Im(\tau)>0\}$. We will also use the shorthand notation for the Dedekind eta function:
$\eta_{n}(\tau):=\eta(n\tau),\qquad n\in\mathbb{Z},\;\tau\in\mathbb{H}.$
@@
Let $f=\sum_{n\in\mathbb{Z}}a_{n}q^{n}$, $f\neq 0$, be such that $a_{n}=0$ for almost all $n<0$. We define the order of $f$ (with respect to $q$) as the smallest integer $N$ such that $a_{N}\neq 0$, and write $N=\operatorname{ord}_{q}(f)$. For instance, let $F=f\circ t=\sum_{n\in\mathbb{Z}}a_{n}t^{n}$ where $t=\sum_{n\geq 1}b_{n}q^{n}$; then the $t$-order of $F$ is defined to be the smallest integer $N$ such that $a_{N}\neq 0$, and we write $N=\operatorname{ord}_{t}(F)$.
@@
For $f:\mathbb{H}\rightarrow\mathbb{C}$ and $m\in\mathbb{N}^{\ast}$, we define $U_{m}(f):\mathbb{H}\rightarrow\mathbb{C}$ by
$U_{m}(f)(\tau):=\frac{1}{m}\sum_{\lambda=0}^{m-1}f\!\left(\frac{\tau+\lambda}{m}\right),\qquad\tau\in\mathbb{H}.$
The operator $U_{m}$ is linear over $\mathbb{C}$; moreover, it is easy to verify that
$U_{mn}=U_{m}\circ U_{n}=U_{n}\circ U_{m},\qquad m,n\in\mathbb{N}^{\ast}.$
The operators $U_{m}$, introduced by Atkin and Lehner [4], are closely related to Hecke operators. They typically arise in the context of partition congruences [1], mostly due to the following property: if
$f(\tau)=\sum_{n=-\infty}^{\infty}f_{n}q^{n}\qquad(q=e^{2\pi i\tau}),$
@@
For $f:\mathbb{H}\rightarrow\mathbb{C}$ we define $U^{(0)}(f),U^{(1)}(f):\mathbb{H}\rightarrow\mathbb{C}$ by $U^{(0)}(f):=U_{2}(Af)$ and $U^{(1)}(f):=U_{2}(Bf)$, where
$A:=\frac{\eta_{8}^{4}\eta_{2}^{2}}{\eta_{4}^{2}\eta_{1}^{4}},\qquad B:=\frac{\eta_{1}^{4}\eta_{2}^{2}\eta_{8}^{4}}{\eta_{4}^{10}}.$
@@
$L_{2\alpha-1}:=U^{(0)}(L_{2\alpha-2}),\qquad L_{2\alpha}:=U^{(1)}(L_{2\alpha-1}).$
The proof of the following lemma is completely analogous to [3, p. 23 ] and is therefore omitted.
@@
$t$ $:=\frac{\eta_{28}\eta_{4}}{\eta_{7}\eta_{1}},~~~~~~~~~~~~~~~~~~~~~p:=\frac{1}{2}\left(1+\frac{7\eta_{14}^{7}\eta_{4}^{5}\eta_{1}^{6}}{\eta_{28}^{3}\eta_{7}^{2}\eta_{2}^{13}}\right),$
$p_{0}$ $:=\frac{\eta_{7}\eta_{4}^{7}}{\eta_{28}\eta_{1}^{7}},~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~y_{0}:=\frac{7\eta_{14}^{5}\eta_{4}^{5}\eta_{2}}{\eta_{28}^{3}\eta_{1}^{8}},~~~~~~~~~~~~~~~~~p_{1}:=\frac{1}{4}(p+2pt+t)+t^{2}.$
@@ 3. Modular equation
Our proof of Radu and Sellers’ conjecture relies on the identities in the Appendix. All these identities can be proved using modular function theory.
Regarding the valence formula in modular forms, Garvan has written a MAPLE package called ETA, which can prove Group I, Group III, and Group IV in the Appendix automatically. See
@@
A tutorial for this package can be found in [7].
The identities in Group II of the Appendix cannot be proved directly, since $p_{1}$ is not an eta-quotient. However, we can use the ETA MAPLE package to prove that
$U^{(0)}(t^{3})$ $=p_{0}(256t^{7}+640t^{6}+864t^{5}+656t^{4}+276t^{3}+64t^{2}+7t)$
(3.2) $\quad-y_{0}(32t^{5}+48t^{4}+28t^{3}+8t^{2}+t);$
$U^{(0)}(t^{4})$ $=p_{0}(1024t^{9}+3072t^{8}+4864t^{7}+4672t^{6}+2832t^{5}+1104t^{4}+260t^{3}+28t^{2})$
(3.3) $\quad-y_{0}(128t^{7}+256t^{6}+224t^{5}+112t^{4}+32t^{3}+4t^{2});$
@@
where y:=2p1y:=2p-1 is an eta-quotient. Then we deduce the identities in Group II using (3)–(3.6) along with the identities in Group I, since
$p_{1}=\frac{1}{8}+\frac{1}{2}t+t^{2}+\frac{1}{8}y+\frac{1}{4}yt.$
@@
With $t=t(\tau)$ as in Definition 2.5, we define:
$a_{0}(t)=-2t^{2}-t,\qquad a_{1}(t)=-4t^{2}-2t.$
@@
(3.7) $a_{j}(t)=\sum_{l=1}^{2}s(j,l)\,2^{\left\lfloor\frac{2l+j-1}{2}\right\rfloor}t^{l}.$
@@
$t_{\lambda}(\tau):=t\!\left(\frac{\tau+\lambda}{2}\right),\qquad\tau\in\mathbb{H}.$
Then, in the polynomial ring $\mathbb{C}(t)[X]$,
@@
$a_{0}(t)$ $=t_{0}t_{1}$
$=-\prod_{\lambda=0}^{1}q^{1/2}\prod_{n=1}^{\infty}\frac{(1-q^{14n})(1-q^{2n})}{(1-(-1)^{\lambda n}q^{7n/2})(1-(-1)^{\lambda n}q^{n/2})}$
$=-q\prod_{n=1}^{\infty}\frac{(1-q^{14n})^{3}(1-q^{2n})^{3}}{(1-q^{7n})^{3}(1-q^{n})^{3}}=-2t^{2}-t,$
where the last equality was verified by Garvan’s MAPLE package ETA (3.1).
@@
where the last equality was again verified by the ETA MAPLE package. ∎
@@
For $u:\mathbb{H}\rightarrow\mathbb{C}$ and $j\in\mathbb{Z}$,
@@
Multiplying both sides by $u_{\lambda}t_{\lambda}^{j-2}$, where $u_{\lambda}(\tau):=u\!\left(\frac{\tau+\lambda}{2}\right)$, gives
@@
Summing both sides over all $\lambda\in\{0,1\}$ completes the proof of the lemma. ∎
According to Lemma 3.3, we can calculate $U^{(0)}(t^{k})$, $U^{(0)}(p_{1}t^{k})$, $U^{(1)}(y_{0}t^{k})$, and $U^{(1)}(p_{0}t^{k})$ using the identities in the Appendix for all $k\geq 1$ recursively. For example,
$U^{(1)}(y_{0}t^{3})$ $=(4t^{2}+2t)U^{(1)}(y_{0}t^{2})+(2t^{2}+t)U^{(1)}(y_{0}t)$
$=p_{1}\bigl(2^{17}t^{13}+2^{19}t^{12}+125\cdot 2^{13}t^{11}+315\cdot 2^{12}t^{10}+2261\cdot 2^{9}t^{9}+1505\cdot 2^{9}t^{8}$
$\qquad+6031\cdot 2^{6}t^{7}+1131\cdot 2^{7}t^{6}+4953\cdot 2^{3}t^{5}+1875\cdot 2^{2}t^{4}+441\cdot 2t^{3}+49t^{2}\bigr)$
$\quad-\bigl(2^{17}t^{15}+9\cdot 2^{16}t^{14}+145\cdot 2^{13}t^{13}+93\cdot 2^{14}t^{12}+2693\cdot 2^{9}t^{11}+3591\cdot 2^{8}t^{10}$
(3.8) $\qquad+7165\cdot 2^{6}t^{9}+5305\cdot 2^{5}t^{8}+5675\cdot 2^{3}t^{7}+1035\cdot 2^{3}t^{6}+463\cdot 2t^{5}+13\cdot 2^{2}t^{4}\bigr).$